# TD-Paint: Faster Diffusion Inpainting Through Time-aware Pixel Conditioning

### v1

@@ 5 Results and Discussion
This section empirically shows that TD-Paint: (a) produces high-quality inpainting with a clean condition on various mask sizes and shapes, on par or better than other inpainting models; (b) provides more efficient sampling, making it faster than other diffusion-based models without requiring a dedicated architecture; (c) generates diverse, high-quality images. Details on the training masks and their corresponding ablation study are presented in Appendix A, and additional qualitative results are provided in Appendix C.
Figure 4: Qualitative results: TD-Paint against state-of-the-art inpainting methods on CelebA-HQ. Zoom in for better details. Additional examples can be found in Appendix C.
Figure 5: Qualitative results: TD-Paint against state-of-the-art inpainting methods on ImageNet1K. Additional examples can be found in Appendix C.
Figure 6: Qualitative results: TD-Paint against state-of-the-art inpainting methods on Places2. Additional examples can be found in Appendix C.
@@ 5.1 Experimental Methodology
Baselines: We evaluate TD-Paint against state-of-the-art diffusion-based inpainting methods: RePaint [1], which conditions the generative process using noisy inputs and synchronizes them with the output through resampling. MCG [4], which adds a correction term to keep the generation closer to the data manifold, and CoPaint [29] that utilizes Tweedie’s formula for better generation. Recent studies [4, 29] have shown that CoPaint and MCG are among the best-performing inpainting methods. We denote the RePaint-20 model as using 20 resampling steps, while RePaint-1 refers to the model with 1 resampling step, matching the number of steps used by TD-Paint. Additionally, TD-Paint is compared with state-of-the-art mask and image conditioned inpainting models: the CNN-based LaMa [16] and transformer-based MAT [17].
Implementation Details: Our approach is validated using the CelebA-HQ [30] dataset, the ImageNet1K [31] dataset, and the Places2 dataset [32] at 256x256 resolution. We modify the implementation of [2], maintaining all their hyperparameters. Training on CelebA-HQ is conducted for approximately 150K steps with batch size 64 on 4 A100, for ImageNet1K and Places2 for about 200K steps with batch size 128 on 8 A100. For baselines, we utilize existing code and pre-trained models when available. For ImageNet1K, we train LaMa for 1M steps on batch size 5 using their implementation, and MAT for 300K steps on batch size 32 using their implementation.
Evaluation Metrics: Image quality is evaluated using established metrics from the inpainting literature: the Learned Perceptual Image Patch Similarity [33] (LPIPS), the Structural Similarity Index Measure [34] (SSIM), and the Kernel Inception Distance [35] (KID) (using the TorchMetrics [36] implementation). The number of diffusion steps (NBS) and the mean time to inpaint an image (Runtime) are used to evaluate the computational efficiency of TD-Paint. For evaluation, we use 2,824 images from the CelebA-HQ test set, 5,000 images from ImageNet1K, and 2,000 images from Places2.
@@ 5.2 Comparison with Diffusion-Based Models
Wide and Narrow masks In the standard image inpainting scenario, TD-Paint is compared using Wide and Narrow masks following LaMa [16] protocol. Tables 1 and 3 shows that TD-Paint consistently outperforms other diffusion-based models, improving by 20% RePaint-20 LPIPS’s on the Wide mask on CelebA-HQ, ImageNet1K and Places2, and by 30% on the Narrow masks. MCG lacks global consistency, resulting in significant artifacts on Wide masks, as seen in Figures 4 and 5 where it produces eyes of different colors, strange textures, and inpainting artifacts. RePaint produces high-quality images at the cost of significant inference time. Our approach can produce high-quality images while requiring much less processing time.
@@
Super-Resolve 2x and Altern. Lines masks In the Super-Resolve 2x setting, every other pixel is removed from the image, while the Altern. Lines setting removes every other line. All diffusion-based baselines achieve low LPIPS scores and produce high-quality images for both types of masks (see Figures 4, 5 and 6). TD-Paint outperforms all considered baselines on CelebA-HQ and comes a close second to CoPaint on ImageNet1K and Places2, as shown in Tables 1 and 3. The third-best performing model is RePaint. Compared to RePaint, TD-Paint improves LPIPS by 115% in the Super-Resolve 2x setting and by 69% in the Altern. Lines setting.
@@
Half and Expand masks The Half setting removes the right part of the images, and the Expand setting keeps only the central 64$\times$64 part of a 256$\times$256 pixel image. Both LPIPS and SSIM metrics are less suitable when a significant portion of the image is missing, as they rely on a single ground truth. This penalizes methods that generate realistic images with semantics different from the ground truth [1]. In this case, the KID, which measures the distance between distributions, is more reliable for assessing image quality. Applying Half and Expand masks is particularly challenging, as they remove substantial portions of the images. task where an important part of the images is removed. This complexity is shown both visually in Figures 4, 5 and 6 and by the quantitative results in Tables 1 and 3. Our model achieves the best performance on both the Half and Expand masks across all datasets, as measured by the KID, while significantly reducing the inference time. Figure 4 shows that TD-Paint can produce high-quality images in this challenging setting where RePaint lacks global consistency. On ImageNet1K and Places2 (see Figures 5 and 6), we observe that MCG and CoPaint often mirror the images from the Half mask, resulting in symmetrical outputs. Similar behavior is observed on Expand masks, with significant texture blending.
@@
Table 1: Quantitative results: LPIPS and SSIM evaluation of diffusion models for inpainting on the CelebA-HQ and ImageNet1K datasets.
Table 2: Quantitative results: evaluation of CNN- and transformer-based models for inpainting on the CelebA-HQ and ImageNet1K datasets.
@@
Table 3: Quantitative results: KID evaluation of diffusion models for inpainting on the CelebA-HQ and ImageNet1K datasets.
5.3 Comparison with CNN- and Transformer-Based Models
Figure 7: Examples of diverse generations using TD-Paint on the CelebA and ImageNet, with the same input image and different initial noise. Additional examples are available in Appendix C.
We compare TD-Paint with LaMa (CNN-based) and MAT (transformer-based) models in Table 2.
Super-Resolve 2x and Altern. Lines masks Table 2 shows that TD-Paint outperforms the baselines by a wide margin. Particularly MAT struggles with this task and often produces images with significant artifacts and blurring (see Figures 4, 5 and 6). In contrast, TD-Paint consistently achieves lower LPIPS scores and higher SSIM values, reflecting superior image quality and performance.
Half and Expand masks On CelebA-HQ, TD-Paint achieves the best results across all datasets, as indicated by the KID metrics (see Table 3). As shown in Figures 4, 5 and 6, LaMa generates blurry artifacts in both the Half and Expand settings, whereas our proposed model consistently produces high-quality images. This behavior of LaMa may be due to overfitting to the training mask distribution, as suggested by previous studies [1].
Table 4: Inpainting speed for different methods.
@@ 5.3 Comparison with CNN- and Transformer-Based Models
We compare the time efficiency of different approaches by computing the average time to sample 100 images consecutively on a single V100 and the results are reported in Table 4. State-of-the-art approaches require over 6×6$\times$6 × longer to sample compared to TD-Paint. The increased time in RePaint is due to the resampling mechanism needed to synchronize condition and generation, while MCG requires an additional backward pass and more steps to optimize the image. In contrast, our model reduces inference time by trading fine-tuning costs, enabling faster generation of high-quality images than other diffusion-based inpainting models.
@@ 5.5 Diversity of Generated Images
While TD-Paint performs fast and high-quality inpainting, we must ask whether this comes at the cost of diversity. To evaluate this, we compute the Diversity Score [37] by generating 10 different images for 100 inputs. The quantitative results reported in Table 5. The most diverse model is RePaint-1, which also has a high LPIPS score. In contrast, TD-Paint achieves a high Diversity Score across most masks while consistently producing high-quality images (see Figure 7) with lower LPIPS scores.
@@ Acknowledgments
This work was financially supported by the ANR Labcom LLisa ANR-20-LCV1-0009. We thank our colleagues at CRIANN, who provided us with the computation resources necessary for our experiments. This work was performed using HPC resources from GENCI–IDRIS (Grant 2024-AD011013862R1). We thank the Natural Sciences and Engineering Research Council of Canada (NSERC) and the Digital Research Alliance of Canada.
@@ Appendix B Image Quality and Diffusion Steps
Appendix C Additional Qualitative Results
We provide additional qualitative inpainting results compared to the state-of-the-art models described in Section 5.
@@ Appendix C Additional Qualitative Results
Additional diversity results on CelebA-HQ and ImageNet1K can be found in Figure 19, and for ImageNet1K with different conditional classes in Figure 20. Through class conditioning, TD-Paint can generate diverse images based on different target classes. This mechanism allows TD-Paint to guide the inpainting process toward specific semantic categories, resulting in more controlled and contextually relevant image completions. This feature is especially useful when the inpainted region must match a particular class, increasing TD-Paint’s flexibility and effectiveness across a range of inpainting tasks.

### v2

@@ 5 Results and Discussion
This section empirically shows that TD-Paint: (a) produces high-quality inpainting with a clean condition on various mask sizes and shapes, on par or better than other inpainting models; (b) provides more efficient sampling, making it faster than other diffusion-based models without requiring a dedicated architecture; (c) generates diverse, high-quality images. Details on the training masks and their corresponding ablation study are presented in Appendix A, and additional qualitative results are provided in Appendix D. In Appendix C, TD-Paint is compared with state-of-the-art mask and image-conditioned inpainting models: the CNN-based LaMa (Suvorov et al., 2022) and transformer-based MAT (Li et al., 2022).
Figure 4: Qualitative results: TD-Paint against state-of-the-art inpainting methods on CelebA-HQ. Zoom in for better details. Additional examples can be found in Appendix D.
Figure 5: Qualitative results: TD-Paint against state-of-the-art inpainting methods on ImageNet1K. Additional examples can be found in Appendix D.
Figure 6: Qualitative results: TD-Paint against state-of-the-art inpainting methods on Places2. Additional examples can be found in Appendix D.
@@ 5.1 Experimental Methodology
Baselines: We evaluate TD-Paint against state-of-the-art diffusion-based inpainting methods in the pixel space: RePaint (Lugmayr et al., 2022), which conditions the generative process using noisy inputs and synchronizes them with the output through resampling. MCG (Chung et al., 2022), which adds a correction term to keep the generation closer to the data manifold, and CoPaint (Zhang et al., 2023a) that utilizes Tweedie’s formula for better generation. Recent studies (Chung et al., 2022; Zhang et al., 2023a) have shown that CoPaint and MCG are among the best-performing inpainting methods. We denote the RePaint-20 model as using 20 resampling steps, while RePaint-1 refers to the model with 1 resampling step matching the number of steps used by TD-Paint. Furthermore, we conduct comparisons with latent diffusion models that have access to the complete context: LDM (Rombach et al., 2022) and ControlNet (Zhang et al., 2023b). Additionally, we compare against foundation model-based latent diffusion approaches, including Blended Latent Diffusion (BLD) (Avrahami et al., 2023), Uni-paint (Yang et al., 2023), and PowerPaint (Zhuang et al., 2025).
Implementation Details: Our approach is validated using the CelebA-HQ (Karras et al., 2018) dataset, the ImageNet1K (Russakovsky et al., 2015) dataset, and the Places2 dataset (Zhou et al., 2018) at 256x256 resolution. We modify the implementation of (Dhariwal & Nichol, 2021), maintaining all their hyperparameters. Training on CelebA-HQ is conducted for approximately 150K steps with batch size 64 on 4 A100, for ImageNet1K and Places2 for about 200K steps with batch size 128 on 8 A100. For baselines, we utilize existing code and pre-trained models when available. For ImageNet1K, we train LaMa for 1M steps on batch size 5 using their implementation, and MAT for 300K steps on batch size 32 using their implementation. Both LDM and ControlNet are trained using computational resources equivalent to TD-Paint. For LDM, the encoded masked image and the downsampled mask provide additional context during the sampling process. For ControlNet, the encoded masked image alone provides additional context.
Evaluation Metrics: Image quality is evaluated using established metrics from the inpainting literature: the Learned Perceptual Image Patch Similarity (Zhang et al., 2018) (LPIPS), the Structural Similarity Index Measure (Wang et al., 2004) (SSIM), and the Kernel Inception Distance (Bińkowski et al., 2018) (KID) (using the TorchMetrics (Nicki Skafte Detlefsen et al., 2022) implementation). The number of diffusion steps (NBS) and the mean time to inpaint an image (Runtime) are used to evaluate the computational efficiency of TD-Paint. For evaluation, we use 2,824 images from the CelebA-HQ test set, 5,000 images from ImageNet1K, and 2,000 images from Places2.
@@ 5.2 Comparison with Diffusion-Based Models
Wide and Narrow masks In the standard image inpainting scenario, TD-Paint is compared using Wide and Narrow masks following LaMa (Suvorov et al., 2022) protocol. Tables 1 and 2 shows that TD-Paint consistently outperforms other diffusion-based models, improving by 20% RePaint-20 LPIPS’s on the Wide mask on CelebA-HQ, ImageNet1K and Places2, and by 30% on the Narrow masks. MCG lacks global consistency, resulting in significant artifacts on Wide masks, as seen in Figures 4 and 5 where it produces eyes of different colors, strange textures, and inpainting artifacts. RePaint produces high-quality images at the cost of significant inference time. Our approach can produce high-quality images while requiring much less processing time. When processing small masks, LDM occasionally produces minor artifacts, as evident in the top rows of Figures 4 and 5. These artifacts are more pronounced in ControlNet, which lacks mask information, resulting in inconsistencies between known and unknown regions. Among foundation-based models, BLD demonstrates superior performance across all datasets, while Uni-paint and PowerPaint show inferior results on CelebA-HQ but achieving better performance on ImageNet1K and Places2. This performance variation may be attributed to BLD’s use of the provided LDM baselines (our early experiments with Stable Diffusion yielded less favorable results), whereas both Uni-paint and PowerPaint utilize Stable Diffusion as their foundation.
@@
Super-Resolve 2x and Altern. Lines masks In the Super-Resolve 2x setting, every other pixel is removed from the image, while the Altern. Lines setting removes every other line. All pixel-based baselines achieve low LPIPS scores and produce high-quality images for both types of masks (see Figures 4, 5 and 6). In contrast, latent-based models prove inadequate for this task due to their downsampling of inpainting masks, which results in critical information loss. TD-Paint outperforms all considered baselines on CelebA-HQ and comes a close second to CoPaint on ImageNet1K and Places2, as shown in Tables 1 and 2. The third-best performing model is RePaint. Compared to RePaint, TD-Paint improves LPIPS by 115% in the Super-Resolve 2x setting and by 69% in the Altern. Lines setting.
@@
Half and Expand masks The Half setting removes the right part of the images, and the Expand setting keeps only the central 64$\times$64 part of a 256$\times$256 pixel image. Both LPIPS and SSIM metrics are less suitable when a significant portion of the image is missing, as they rely on a single ground truth. This penalizes methods that generate realistic images with semantics different from the ground truth (Lugmayr et al., 2022). In this case, the KID, which measures the distance between distributions, is more reliable for assessing image quality. Applying Half and Expand masks is particularly challenging, as they remove substantial portions of the images. task where an important part of the images is removed. This complexity is shown both visually in Figures 4, 5 and 6 and by the quantitative results in Tables 1 and 2. Our model performs best on both the Half and Expand masks across all datasets (except for CelebA-HQ where it comes close to LDM), as measured by the KID, while significantly reducing the inference time. Figure 4 shows that TD-Paint can produce high-quality images in this challenging setting where RePaint lacks global consistency. On ImageNet1K and Places2 (see Figures 5 and 6), we observe that MCG and CoPaint often mirror the images from the Half mask, resulting in symmetrical outputs. Similar behavior is observed on Expand masks, with significant texture blending. ControlNet exhibits limited generalization capability when handling very large masks, likely due to the absence of mask information, whereas LDM demonstrates robust performance in these scenarios. Our analysis reveals that foundation-based models yield inferior qualitative results on CelebA-HQ and Places2 datasets. However, they show improved performance on ImageNet1K, where class names provide supplementary textual information, which is particularly beneficial in scenarios with limited contextual information.
@@
Table 1: Quantitative results: LPIPS and SSIM evaluation of diffusion models for inpainting on the CelebA-HQ, ImageNet1K and Places2 datasets.
@@
Table 2: Quantitative results: KID evaluation of diffusion models for inpainting on the CelebA-HQ, ImageNet1K and Places2 datasets.
@@
Table 3: Inpainting speed for different diffusion model in the pixel space.
We compare the time efficiency of different diffusion approaches working in the pixel space by computing the average time to sample 100 images consecutively on a single V100, and the results are reported in Table 3. State-of-the-art approaches require over 6×6$\times$6 × longer to sample compared to TD-Paint. The increased time in RePaint is due to the resampling mechanism needed to synchronize condition and generation, while MCG requires an additional backward pass and more steps to optimize the image. In contrast, our model reduces inference time by trading fine-tuning costs, enabling faster generation of high-quality images than other diffusion-based inpainting models.
Figure 7: Examples of diverse generations using TD-Paint on the CelebA and ImageNet, with the same input image and different initial noise. Additional examples are available in Appendix D.
@@ 5.4 Diversity of Generated Images
While TD-Paint performs fast and high-quality inpainting, we must ask whether this comes at the cost of diversity. To evaluate this, we compute the Diversity Score (Lugmayr et al., 2021) by generating 10 different images for 100 inputs. The quantitative results reported in Table 4. The most diverse model is RePaint-1, which also has a high LPIPS score. In contrast, TD-Paint achieves a high Diversity Score across most masks while consistently producing high-quality images (see Figure 7) with lower LPIPS scores.
@@ Acknowledgments
This work was financially supported by the ANR Labcom LLisa ANR-20-LCV1-0009. We thank CRIANN, who provided us with the computation resources necessary for our experiments. This work was performed using HPC resources from GENCI–IDRIS (Grant 2024-AD011013862R1). We thank the Natural Sciences and Engineering Research Council of Canada (NSERC) and the Digital Research Alliance of Canada.
@@ Appendix B Image Quality and Diffusion Steps
Appendix C Comparison with CNN- and Transformer-Based Models
We compare TD-Paint with LaMa (CNN-based) and MAT (transformer-based) models in Tables 6 and 7.
Super-Resolve 2x and Altern. Lines masks Table 6 shows that TD-Paint outperforms the baselines by a wide margin. Particularly MAT struggles with this task and often produces images with significant artifacts and blurring (see Figures 10, 11 and 12). In contrast, TD-Paint consistently achieves lower LPIPS scores and higher SSIM values, reflecting superior image quality and performance.
Half and Expand masks On CelebA-HQ, TD-Paint achieves the best results across all datasets, as indicated by the KID metrics (see Table 7). As shown in Figures 10, 11 and 12, LaMa generates blurry artifacts in both the Half and Expand settings, whereas our proposed model consistently produces high-quality images. This behavior of LaMa may be due to overfitting to the training mask distribution, as suggested by previous studies (Lugmayr et al., 2022).
Table 6: Quantitative results: evaluation of CNN- and transformer-based models for inpainting on the CelebA-HQ, ImageNet1K and Places2 datasets.
Table 7: Quantitative results: KID evaluation of CNN- and transformer-based models for inpainting on the CelebA-HQ, ImageNet1K and Places2 datasets.
Figure 10: Qualitative results: TD-Paint against state-of-the-art inpainting CNN- and transformer-based models on CelebA-HQ. Zoom in for better details. Additional examples can be found in Appendix D.
Figure 11: Qualitative results: TD-Paint against state-of-the-art inpainting CNN- and transformer-based models on ImageNet1K. Additional examples can be found in Appendix D.
Figure 12: Qualitative results: TD-Paint against state-of-the-art inpainting CNN- and transformer-based models on Places2. Additional examples can be found in Appendix D.
Appendix D Additional Qualitative Results
We provide additional qualitative inpainting results compared to the state-of-the-art models described in Sections 5 and C.
@@ Appendix D Additional Qualitative Results
Additional diversity results in CelebA-HQ and ImageNet1K can be found in Figure 22, and for ImageNet1K with different conditional classes in Figure 23. Through class conditioning, TD-Paint can generate diverse images based on different target classes. This mechanism allows TD-Paint guiding the inpainting process toward specific semantic categories, resulting in more controlled and contextually relevant image completions. This feature is especially useful when the inpainted region must match a particular class, increasing TD-Paint’s flexibility and effectiveness across a range of inpainting tasks.
Additional examples of inpainting using object-focused and region-specific masks are presented in Figures 24 and 25. These examples feature user-drawn masks that naturally follow object boundaries and regional structures, demonstrating TD-Paint’s effectiveness in practical image manipulation scenarios. Such realistic mask shapes better reflect how users interact with inpainting tools in real-world applications, where selections typically correspond to meaningful objects or regions rather than arbitrary geometric patterns.
@@
Figure 24: Demonstration of TD-paint application on the ImageNet1K dataset. The figure shows user-drawn masks highlighting specific regions or objects, followed by four generated image variations for each mask.
Figure 25: Demonstration of TD-paint application on the Places2 dataset. The figure shows user-drawn masks highlighting specific regions or objects, followed by four generated image variations for each mask.